OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. x + 2y = 12 y = x + 3 A. (5, 8) B. (2, 5) C. (2, 1) D. (5, 2)

OpenStudy (anonymous):

You can substitute y or x. I suggest you pick y because the second equation is already solved for y All you have to do is isolate the y in the first equation. Can you do that?

OpenStudy (anonymous):

tHAT WOULD GIVE ME B RIGHT? OR WRONG?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Did you just guess?

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

oh ok well good job then.

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow x + 2(x + 3) = 12 }\) I replaced y with x + 3 as we already know y = x + 3

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow x + 2x + 6 = 12 }\) \(\Large \color{purple}{\rightarrow 3x + 6 = 12 }\) \(\Large \color{purple}{\rightarrow 3x = 6 }\) \(\Large \color{purple}{\rightarrow x = 2 }\)

Parth (parthkohli):

http://openstudy.com/users/parthkohli#/updates/4fa7aa68e4b029e9dc388281 See this tutorial for more.